Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Siblings

Expert replies
by gmatblood » Thu Nov 03, 2011 11:18 am
In a room filled with 7 people, 4 people have exactly 1 sibling in the room and 3 people have exactly 2 siblings in the room. If two individuals are selected from the room at random, what is the probability that those two individuals are NOT siblings?

5/21

3/7

4/7

5/7

16/21
Join the discussion
Source: — Problem Solving |

by shankar.ashwin » Thu Nov 03, 2011 11:44 am
P(Not sibling) = 1 - P(sibling)

Let the siblings be (AB) (CD) (EFG)

Picking up AB - 1 possibility
Picking up CD - 1 possibility
Picking up 2 from EFG - 3C2 - 3 possibilities

Total = 5.

Total Number of ways of picking up from 7 - 7C2 = 21.

1 - 5/21 = 16/21. E IMO
Join the discussion

by CappyAA » Thu Nov 03, 2011 11:50 am
The wording can throw you off, but basically I picture this group of 7 as follows: There are 7 people (A, B, C, D, E, F, and G) split into 2 groups. 1 group of 4 people has exactly 1 sibling in the room. This means that there are 2 pairs of siblings. We will assume A & B are siblings and C & D are siblings. The secong group of 3 people has exactly 2 siblings in the group - each other. So E, F, and G are siblings with each other. We want to find the probability that two people selected are NOT siblings.

First we can find the total number of ways to pick two people out of a group of 7. Since the order does not matter, it is 7!/(5!*2!) or 21 different ways. There is 1 way that each of the 2 pairs of siblings can be selected and 3 ways that the trio of siblings can be selected. This is a total of 5 ways siblings can be selected.

Since this is a very small group, we can list the ways that siblings can be selected:

- A&B
- C&D
- E&F
- F&G
- E&G

There are 5 ways siblings can be selected so there must be 21-5 = 16 ways where siblings are not selected. So the answer is 16/21 or E.
Taking the GMAT Again...PhD this time!

October 2008 Score: GMAT - 750 (50 Q, 41 V) :D

Manhattan GMAT 1 - 11/20/11 - 750 (50 Q, 42 V)
Manhattan GMAT 2 - 12/3/11 - 780 (51 Q, 45 V)
Join the discussion

by gmatblood » Thu Nov 03, 2011 2:58 pm
Its all about rephrasing the question!! Thanks a lot CappyAA :)
Join the discussion